OpenStudy (anonymous):

A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground?

OpenStudy (anonymous):

You have to use the ratio for length and measure of a circle to find it

OpenStudy (anonymous):

firstly you need to calculate the circumference of the wheel. to do this, take C = Circumference. So: \[C = 2*\pi*r\] where r = radius of the wheel which is 26". So thus C = 163.36" The full circumference of the wheel is 360 degrees, but you only want 80 degrees. So the length that touched the ground = 163.36*80/360 = 36.302"

OpenStudy (anonymous):

I forgot to convert something :/

OpenStudy (amistre64):

you need angle*r ; but angle has to be in radians

OpenStudy (amistre64):

80 needs a pi so; 80pi/180 is the conversion

OpenStudy (anonymous):

I got 11.56 inches the first time

OpenStudy (amistre64):

26(80)pi/180 .... simplify as wanted

OpenStudy (anonymous):

amistre....that doesn't help me.....I already know that

OpenStudy (anonymous):

I wasn't asking for help

OpenStudy (anonymous):

you dont need to convert it to radians..its not necessary

OpenStudy (amistre64):

i wasnt answering any question you posted :)

OpenStudy (amistre64):

of course you need to convert it to radians; or at least its more helpful that way

OpenStudy (anonymous):

It was an input mistake on my calculator ......11.56*pi = 36.30 inches.....somehow that's where I made the mistake

OpenStudy (anonymous):

hey can u come on google voice i wanna talk to u

OpenStudy (anonymous):

Cherri, how are you going to ask me to come somewhere when you're not there?